OpenStudy (anonymous):

i cant get these two questions please help |1+p/2| = 3 4,8 -4,-8 -4,8 4,-8 |3r - 12| = 0 -4,4 0,4 0,-4 4

OpenStudy (mertsj):

The first equation means that 1+p/2 is exactly 3 units from 0 on the number line. That means it could be either 3 or -3. So write and solve the two equations: 1+p/2 = 3 or 1+p/2 = -3

OpenStudy (anonymous):

that didnt help but thanks i guess

OpenStudy (mertsj):

What. You don't know how to solve the two equations?

OpenStudy (anonymous):

if I did I wouldnt have came here.

OpenStudy (mertsj):

To solve 1+p/2 = 3 first subtract 1 from both sides and get p/2=2. Now multiply both sides by 2 and get p = 4.

OpenStudy (mertsj):

To solve 1+p/2 = -3 first subtract 1 from both sides and get p/2 = -4. Now multiply both sides by 2 and get p = -8

OpenStudy (mertsj):

|3r-12|=0 can only be true if 3r-12 = 0 Add 12 to both sides and get 3r = 12. Now divide both sides by 3 and get r = 4